### v1

# A Dimca-Greuel type inequality for foliations

Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. We prove the following Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}[[x,y]]$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel’s conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
Key words and phrases:
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
2020 Mathematics Subject Classification:
Primary 32S65 - 32M25
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done (Spanish grant PID2019-105896GB-I00 funded). This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979.
1. Introduction
Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables. Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$. The Milnor number of $C$ is
$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
and its Tjurina number is
$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
Is clear from their definitions that $\tau(C)\leq\mu(C)$.
In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [10], and [15] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. Moreover if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9], [11]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$. If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
If the foliation $\mathcal{F}$ is of the second type and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$ then we get a bound of the Dimca-Greuel type for foliations.
Our main theorem is
Theorem A.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
$\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
We will prove Theorem A in Section 3.
As a consequence of Theorem A we deduce (see Corollary 3.5) the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15].
2. Multiplicity of a foliation along a divisor of separatrices
Throughout this text $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
(2) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, page 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
(3) $\theta(t)=\begin{cases}-\frac{Q(\gamma(t))}{x^{\prime}(t)}&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{P(\gamma(t))}{y^{\prime}(t)}&\text{if $y(t)\neq 0$}.\end{cases}$
We extend the notion of multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}B$ of separatrices of $\mathcal{F}$ at $p$ in the following way:
(4) $\mu_{p}(\mathcal{F},\mathcal{B})=\left(\displaystyle\sum_{B}a_{B}\mu_{p}(\mathcal{F},B)\right)-\deg(\mathcal{B})+1.$
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
(5) $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we get
(6) $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$. See [12] for details.
Remark 2.1.
Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (2) and (3), we have
(7) $\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$
3. Proof of Theorem A
Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
(8) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (8):
(9) $\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Denote by $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([15]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [15, Claim 4.4] we have
(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},$
for $\nu_{p}(C)\geq 2$.
We will need the following lemma.
Lemma 3.1.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$, then
(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Moreover, the equality holds if and only if $B$ is smooth.
Proof.
Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we get
(12) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
where
(13) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behaviour under blow-up of the GSV index [3, p.30] we get
(14) $GSV_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and by [7, Proposition 6.2] we have
(15) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
After [15, Equation (4)] we get:
(16) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (10). From (15), (14), and (16) we deduce
(17) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}+\alpha.$
By (12), (14) and (17) it holds
(18) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$ $=$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
-- $\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha.$
Hence, if $B$ is a singular curve then the lemma follows by (10).
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some formal power series $h(x,y)\in\mathbb{C}[[x,y]]$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(19) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (15) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality.
Moreover, if (11) is an equality, then $B$ is necessarily smooth. Indeed if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (18) we have that (11) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
Corollary 3.2.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$ after a composition of $\ell$ blows-up then
(20) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Proof.
It follows applying Lemma 3.1 $\ell$ times. ∎
Remark 3.3.
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(21) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (19) if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
Remark 3.4.
If $\mathcal{F}$ is a non-dicritical second type foliation at $p\in\mathbb{C}^{2}$, and $B$ is the only separatrix of $\mathcal{F}$, then $\mathcal{F}$ is a generalized curve foliation. In fact, assume by contradiction that $\mathcal{F}$ admits a saddle-node singularity in its reduction process with strong separatrix on the exceptional divisor. Without lost of generality, we can suppose that $\mathcal{F}$ is a reduced foliation after a blow-up $\sigma$ at $p$. Since $\mathcal{F}$ is of second type, after a blow-up, the strong separatrix of the strict transform $\tilde{\mathcal{F}}$ of $\mathcal{F}$ through $q\in\sigma^{-1}(p)$ is contained in $E=\sigma^{-1}(p)$, thus the Camacho-Sad index is $CS_{q}(\tilde{\mathcal{F}},E)=0$, but it is a contradiction with Camacho-Sad’s formula
$CS_{q}(\tilde{\mathcal{F}},E)=E\cdot E=-1.$
Proof of Theorem A: If $B$ is a non-singular curve then, by Remark 3.3, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$ and the theorem follows.
Let us suppose now that $B$ is a singular separatrix, that is $\nu_{p}(B)\geq 2$. After Corollary 3.2 we can suppose, without lost of generality that the number of blows-up $n$ needed in the process of the reduction of the singularity of $B$ is one. In particular, after a blow-up $\sigma$ the strict transform $\tilde{B}$ of $B$ is smooth, that is the algebraic multiplicity of $\tilde{B}$ is one.
Applying Lemma 3.1 and after (14) we get
(22) $[3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})]-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Since $\tilde{B}$ is smooth, from (22) and the equality $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{q}(\tilde{\mathcal{F}},\tilde{B})$ (see Remark 3.3) we obtain
(23) $-\tau_{q}(\tilde{\mathcal{F}},\tilde{B})-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Moreover, by the definition of $m_{p}(\mathcal{F})$ and after [11, Proposition 3.3], if $\mathcal{B}=B+\mathcal{B}^{\prime}$ is a balanced divisor of separatrices of $\mathcal{F}$ (for some divisor $\mathcal{B}^{\prime}$, if $B$ is the only separatrix of $\mathcal{F}$ we consider for $\mathcal{B}^{\prime}$ the empty divisor) then we have
$m_{p}(\mathcal{F})-\nu_{p}(B)=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{B}^{\prime})-1+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical}\\ \nu_{p}(\mathcal{B}^{\prime})+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is dicritical},\end{array}\right.$
where $\xi_{p}(\mathcal{F})$ is the tangency excess of $\mathcal{F}$ (see for example [7, Definition 2.3]). So $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$ except in the case that $\mathcal{F}$ is a foliation of the second type ($\xi_{p}(\mathcal{F})=0$) and $B$ is its only separatrix ($\nu_{p}(\mathcal{B}^{\prime})=0$) in which case we get $m_{p}(\mathcal{F})-\nu_{p}(B)=-1$. In the case $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$, from (23) we obtain $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. Otherwise, assume that $\mathcal{F}$ is of the second type and that $B$ is its only separatrix. By Remark 3.4, $\mathcal{F}$ is a generalized curve foliation, so $GSV_{p}(\mathcal{F},B)=0$. Moreover, since $\tilde{B}$ is smooth then the left-term of (11) is zero, so the theorem follows by Lemma (3.1) (since $B$ is singular the inequality (11) is strict) and Remark 3.3 ($\tilde{B}$ is smooth). $\Box$
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15]:
Corollary 3.5.
Let $f(x,y)\in\mathbb{C}[[x,y]]$ an irreducible power series. Put $C:f(x,y)=0$ the branch defined by $f$. Then
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
Proof.
It is a consequence of Remark 2.1 and of applying Theorem A to foliation $\mathcal{F}:df=0$. ∎
Remark 3.6.
In [15] Wang proved that the number $3\mu_{p}(B)-4\tau_{p}(B)$ is a non-negative number, non-decreasing under blowups and strictly increasing unless the curve $B$ is non-singular. The result of Wang is an special case of Lemma 3.1, taking $\mathcal{F}=df$ where $B$ is given by $f(x,y)=0$. Lemma 3.1 and Theorem A generalizes this for the number $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$.
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
Corollary 3.7.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ be a balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(24) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
Proof.
By (6) we have $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}$. Hence we will prove that
(25) $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
Suppose that $\mathcal{B}_{0}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
$\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
Thus
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})<-\ell+1=3\ell-4\ell+(4-3)$
and
$3\left(\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-\ell+1\right)<4\left(\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})-\ell+1\right),$
so
$\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
The following example illustrates Corollary 3.7.
Example 3.8.
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$.
Let $C=\{x=0\}$, $\mathcal{B}_{0}=C+C_{0,1}+C_{1,1}$, $\mathcal{B}_{\infty}=C_{2,1}$, then $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices for $\mathcal{F}$. We have
$\mu_{0}(\mathcal{F},\mathcal{B}_{0})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{0,1})+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+29+21-2+1=54,$
$\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=\mu_{0}(\mathcal{F},C_{2,1})=21,$
and
$T_{0}(\mathcal{F},\mathcal{B}_{0})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{0,1})+\tau_{0}(\mathcal{F},C_{1,1})=5+28+21=54.$
Hence
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},\mathcal{B}_{0})-\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})+1=54-21+1=34,$
and
$\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B}_{0})}=\frac{34}{54}<\frac{4}{3}.$
We conclude the paper with the following question:
Question: Is the inequality (24) true if instead of $T_{p}(\mathcal{F},\mathcal{B}_{0})$ we write $\tau_{p}(\mathcal{F},\mathcal{B}_{0})$?