### v2

# A Dimca-Greuel type inequality for foliations

Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. Under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we prove the Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
Key words and phrases:
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
2020 Mathematics Subject Classification:
Primary 32S65 - 32M25
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done (Spanish grant PID2019-105896GB-I00 funded). This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979.
1. Introduction
Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables. Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$. The Milnor number of $C$ is
$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
and its Tjurina number is
$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
Is clear from their definitions that $\tau(C)\leq\mu(C)$.
In [7, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [11], and [19] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. The algebraic multiplicity $\nu_{p}(\mathcal{F})$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$. Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [10], [12]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$. If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
If the foliation $\mathcal{F}$ is of the second type (see [10, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [9, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [15] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [8, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$, then, under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we will obtain a bound of the Dimca-Greuel type for foliations. More specifically, in Foliation theory, two type of separatrices of $\mathcal{F}$ are distinguished: isolated separatrices $\text{ISO}_{p}(\mathcal{F})$ and dicritical separatrices $\text{Dic}_{p}(\mathcal{F})$. Isolated separatrices of $\mathcal{F}$ correspond to separatrices that are images, by the reduction process of singularities $\pi$ of $\mathcal{F}$, of separatrices of non-degenerate singularities, or weak or strong separatrices of saddle-nodes singularities in a reduction of singularities of $\mathcal{F}$. On the other hand, dicritical separatrices correspond to separatrices that are images, by $\pi$, of separatrices that intersect transversally with a dicritical divisor in the last step of the reduction of singularities of $\mathcal{F}$. For more details about these definitions, see Section 3.
In order to establish an inequality of Dimca-Greuel type, it is necessary to introduce a number $\text{DG}_{p}(\mathcal{F},B)$ that depends on the reduction of singularities of the pair $(\mathcal{F},B)$ and on the type of separatrix that $B$ is, see Definition 21. We will see that the negativity of this number is crucial in order to obtain an inequality of the Dimca-Greuel type for foliations.
Our main theorem is
Theorem A.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\text{DG}_{p}(\mathcal{F},B)<0$, then
(2) $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
We will prove Theorem A in Section 4. In Example 4.3, we will show that we cannot removed the hypothesis over $\text{DG}_{p}(\mathcal{F},B)$.
As a consequence of Theorem A we deduce (see Corollary 5.1) the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
2. Multiplicity of a foliation along a divisor of separatrices
Throughout this note $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
(3) $v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve. We denote by $\text{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$ at $p$.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by a vector field $v$ as in (3) and $B\in\text{Sep}_{p}(\mathcal{F})$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
(4) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [5, p. 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
(5) $\theta(t)=\begin{cases}-\frac{Q(\gamma(t))}{x^{\prime}(t)}&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{P(\gamma(t))}{y^{\prime}(t)}&\text{if $y(t)\neq 0$}.\end{cases}$
We extend the notion of multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}B$ of separatrices of $\mathcal{F}$ at $p$ in the following way:
(6) $\mu_{p}(\mathcal{F},\mathcal{B})=\left(\displaystyle\sum_{B}a_{B}\mu_{p}(\mathcal{F},B)\right)-\deg(\mathcal{B})+1.$
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [14, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
(7) $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
(8) $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$. See [13] for details.
Remark 2.1.
Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (4) and (5), we have
(9) $\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [8, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [8, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [8, Proposition 6.2], we obtain
(10) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
Consequently when $B$ is quasihomogeneous by ([17], Satz p.123 a) and d)] and (10) or by [20, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
3. Reduction of singularities of foliations and their separatrices
We say that $p\in\mathbb{C}^{2}$ is a reduced singularity for $\mathcal{F}$ if the linear part $\text{D}v(p)$ of the vector field $v$ in (3) is non-zero and has eigenvalues $\lambda_{1},\lambda_{2}\in\mathbb{C}$ fitting in one of the cases:
  1. (i)
    $\lambda_{1}\lambda_{2}\neq 0$ and $\lambda_{1}\lambda_{2}\not\in\mathbb{Q}^{+}$ (non-degenerate);
  2. (ii)
    $\lambda_{1}\neq 0$ and $\lambda_{2}\neq 0$ (saddle-node singularity).
In the case $(i)$, there is a system of coordinates $(x,y)$ in which $\mathcal{F}$ is defined by the equation
(11) $\omega=x(\lambda_{1}+a(x,y))dy-y(\lambda_{2}+b(x,y))dx,$
where $a(x,y),b(x,y)\in{\mathbb{C}}[[x,y]]$ are non-units, so that $\text{Sep}_{p}(\mathcal{F})$ is formed by two transversal analytic branches given by $\{x=0\}$ and $\{y=0\}$. In the case $(ii)$, up to a formal change of coordinates, the saddle-node singularity is given by a 1-form of the type
(12) $\omega=x^{k+1}dy-y(1+\lambda x^{k})dx,$
where $\lambda\in\mathbb{C}$ and $k\in\mathbb{Z}^{+}$ are invariants after formal changes of coordinates (see [16, Proposition 4.3]). The curve $\{x=0\}$ is an analytic separatrix, called strong separatrix, whereas $\{y=0\}$ corresponds to a possibly formal separatrix, called weak separatrix. The integer $k+1>1$ is called tangency index of $\mathcal{F}$ with respect to the weak separatrix.
Let $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ be a composition of blow-ups maps. The divisor $\mathcal{D}$ is a finite union of components which are embedding projective lines, crossing normally at corners. If $\mathcal{F}$ is the foliation given by the 1-form $\omega$, we denote ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ the strict transform of $\mathcal{F}$, the germ of foliation on $(\tilde{X},\mathcal{D})$ defined locally by $\pi^{*}(\omega)$, after cancelling one-dimensional singular components. For a uniform analysis, we include the possibility of $\pi$ being the identity map and, abusing notation, we set in this case $\tilde{X}=\mathbb{C}^{2}$, $\mathcal{D}=\{p\}$ and $\tilde{\mathcal{F}}=\mathcal{F}$.
With respect to the divisor $\mathcal{D}$, the foliation $\tilde{\mathcal{F}}$ at a point $q\in\mathcal{D}$ can be:
  • Regular, if there are local analytic coordinates $(x,y)$ at $q$ such that $\mathcal{D}\subset\{xy=0\}$ and ~$\tilde{\mathcal{F}}:dx=0$.
  • Singular; if it is not regular.
  • Reduced, if $q$ is a reduced singularity for $\tilde{\mathcal{F}}$ and $\mathcal{D}\subset\text{Sep}_{q}(\tilde{\mathcal{F}})$.
When $\mathcal{D}=\{p\}$, these notions coincide with the ordinary concepts of regular point, singular point and reduced singularity. We say that $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ is a reduction of singularities for $\mathcal{F}$ if all points $q\in\mathcal{D}$ are either regular or reduced singularities. There always exists a reduction of singularities [18]. Besides, there exists a minimal one, in the sense that it factorizes any other reduction of singularities by an additional sequence of blow-ups. All along this note, reductions of singularities are supposed to be minimal.
Given a germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ with a fixed reduction process $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},0)$ for $\mathcal{F}$, a component $D\subset\mathcal{D}$ can be:
  • Non-dicritical, if $D$ is $\tilde{\mathcal{F}}$-invariant. In this case $D$ contains a finite number of reduced singularities. Each non-corner singularity carries a separatrix transversal to $D$, whose projection by $\pi$ is a curve in $\text{Sep}_{p}(\mathcal{F})$.
  • Dicritical, if $D$ is not $\tilde{\mathcal{F}}$-invariant. The definition of reduction of singularities gives that $D$ may intersect only non-dicritical components and that $\tilde{\mathcal{F}}$ is everywhere transverse to $D$. The $\pi$-image of a local leaf of $\mathcal{F}$ at each non-corner point of $D$ belongs to $\text{Sep}_{p}(\mathcal{F})$.
Denote by $\text{Sep}_{p}(D)\subset\text{Sep}_{p}(\mathcal{F})$ the set of separatrices whose strict transforms by $\pi$ intersect the component $D\subset\mathcal{D}$. If $B\in\text{Sep}_{p}(D)$ with $D$ non-dicritical, $B$ is said to be isolated. Otherwise, it is said to be a dicritical separatrix. This engenders the decomposition $\text{Sep}_{p}(\mathcal{F})=\text{ISO}_{p}(\mathcal{F})\cup\text{Dic}_{p}(\mathcal{F})$, where notation is self evident. The set $\text{ISO}_{p}(\mathcal{F})$ is finite and contains all purely formal separatrices. It subdivides further in two classes: weak separatrices – those arising from the weak separatrices of saddle-nodes – and strong separatrices – corresponding to strong separatrices of saddle-nodes and separatrices of non-degenerate singularities. On the other hand, if $\text{Dic}_{p}(\mathcal{F})$ is a non-empty set then it is an infinite set of analytic separatrices of $\mathcal{F}$. A foliation $\mathcal{F}$ is said to be dicritical when $\text{Sep}_{p}(\mathcal{F})$ is infinite, which is equivalent to saying that $\text{Dic}_{p}(\mathcal{F})$ is non-empty. Otherwise, $\mathcal{F}$ is called non-dicritical. When we deal with a simple blow-up $\sigma$ at $p\in\mathbb{C}^{2}$, we say that $\sigma$ is dicritical with respect to $\mathcal{F}$, if the exceptional divisor $\sigma^{-1}(p)$ is not $\tilde{\mathcal{F}}$-invariant. Otherwise, $\sigma$ is called non-dicritical.
Let $\mathcal{F}$ be a holomorphic foliation at $(\mathbb{C}^{2},p)$ and $B\in\text{Sep}_{p}(\mathcal{F})$ On a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we mark the reduction process of singularities of the pair $(\mathcal{F},B)$:
(13) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ is the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(j+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
Let
(14) $\alpha_{j}:=\left\{\begin{array}[]{ll}\dim\frac{(\sigma_{j+2}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j+1)}}}{(\sigma_{j+1}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j)}}},&\hbox{ if $j=0,\ldots,N-2$}\\ \dim\frac{\Omega_{B^{(N)}}}{\sigma_{N}^{*}\Omega_{B^{(N-1)}}},&\hbox{ if $j=N-1$}\end{array}\right.$
where for a curve $B$, $\Omega_{B}$ denotes the module of Kähler differentials on $B$. We note that $\alpha_{0}$ appears as a term in [19, Equation 4]. Additionally, notice that $\alpha_{j}$ is the $\alpha_{0}$ for $B^{(j)}$. Moreover, we observe that $\alpha_{j}\geq 0$, and the equality holds when the curve $B^{(j)}$ is smooth.
We set
(15) $R_{p}(\mathcal{F},B):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
where
(16) $m_{q^{(j)}}(\mathcal{F}^{(j)})=\left\{\begin{array}[]{ll}\nu_{q^{(j)}}(\mathcal{F}^{(j)})+1&\hbox{\rm if $\sigma_{j+1}$ is dicritical}\\ \nu_{q^{(j)}}(\mathcal{F}^{(j)})&\hbox{\rm if $\sigma_{j+1}$ is non-dicritical}.\\ \end{array}\right.$
Notice that for $j=0$ we get $q^{(0)}=p$, $\mathcal{F}^{(0)}=\mathcal{F}$ and the equality (16) for $j=0$ becomes:
(17) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical},\\ \end{array}\right.$
where $\sigma$ is a blow-up at the point $p$.
The following lemma relates $R_{p}(\mathcal{F},B)$ to the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ (respectively $\mathcal{F}^{(N)}$) with respect to $B$ (respectively $B^{(N)}$).
Lemma 3.1.
With the above notations we get
$R_{p}(\mathcal{F},B)=GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right].$
Proof.
From blow-up formula of the GSV index [4, p.30] we get
(18) $GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})$ $=$ $GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})$
$+$ $\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})),\,\,\,\,\hbox{\rm for }\,\,\,0\leq j\leq N-1.$
Up substituting the equality (18) in (15), we get
$R_{p}(\mathcal{F},B)$ $=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}))\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})-GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right],$
and the lemma follows. ∎
For a germ of a holomorphic foliation $\mathcal{F}$ and a separatrix $B\in\text{Sep}_{p}(\mathcal{F})$ at $(\mathbb{C}^{2},p)$, we introduce the number $\text{DG}_{p}(\mathcal{F},B)$ as follows:
(21) $\text{DG}_{p}(\mathcal{F},B)$ ::=:= $\left\{\begin{array}[]{ll}R_{p}(\mathcal{F},B)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})&\hbox{\rm if $B\in\text{ISO}_{p}(\mathcal{F})$}\\ R_{p}(\mathcal{F},B)&\hbox{\rm if $B\in\text{Dic}_{p}(\mathcal{F})$}.\\ \end{array}\right.$
From (14), we deduce that $\alpha_{i}$ are analytic invariants, thus we have $\text{DG}_{p}(\mathcal{F},B)$ is also an analytic invariant depending of $\mathcal{F}$ and $B$. As a consequence of Lemma 3.1, we can establish $\text{DG}_{p}(\mathcal{F},B)$ in terms of the GSV-indices of the foliations $\mathcal{F}$ and $\mathcal{F}^{(N)}$ and some numbers depending of the resolution of the singularities of $B$.
4. Proof of Theorem A
First of all, we will study the behaviour of $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ under a blow-up $\sigma$ at $p\in\mathbb{C}^{2}$. For this, we consider $\alpha_{0}$ as in (14) for $j=0$.
We have the following lemma.
Lemma 4.1.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B\in\text{Sep}_{p}(\mathcal{F})$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by a blow-up $\sigma$ at $p$ then
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0},$
where $q\in\tilde{B}\cap\sigma^{-1}(p)$, and $\alpha_{0}$ is as in (14).
Proof.
By [6, Equality (1.2), p. 291] we get
(22) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
where $m_{p}(\mathcal{F})$ was introduced in (17).
On the other hand by the behaviour under blow-up of the GSV index (see [4, p.30]) we get
(23) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and after [8, Proposition 6.2] we have
(24) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
From [19, Equation (4)] we obtain
(25) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha_{0},$
where $\alpha_{0}$ was defined in (14). From (24), (23), and (25) we deduce
(26) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha_{0}.$
By (22), (23) and (26) it holds
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0}.$
Remark 4.2.
If $\mathcal{F}$ is a reduced foliation at $p\in\mathbb{C}^{2}$, and $B$ is a smooth separatrix through $p$, then, by Remark 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F}.B)=\tau_{p}(\mathcal{F},B)$. Therefore, according to [3, p. 538], we have $\tau_{p}(\mathcal{F},B)=1$, if $p$ is a non-degenerate singularity, or if $p$ is a saddle-node singularity and $B$ is the strong separatrix; and $\tau_{p}(\mathcal{F},B)=k+1$, if $p$ is a saddle-node singularity and $B$ is the weak separatrix. Here $k\in\mathbb{Z}^{+}$ is a formal invariant in the case that $p$ is a saddle-node singularity of $\mathcal{F}$. In any cases, $\tau_{p}(\mathcal{F},B)>0$.
We will now prove the main theorem of this paper.
4.1. Proof of Theorem A
Let $\mathcal{F}$ be a holomorphic foliation on $(\mathbb{C}^{2},p)$, and let $B$ be a separatrix of $\mathcal{F}$ at $p$. Then, on a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we marked the reduction process of the pair $(\mathcal{F},B)$:
(27) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ coincide with the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$. Put $\mu^{(i)}=\mu_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$, for any $i\in\{0,\ldots,N\}$.
Applying Lemma 4.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\overset{\sigma_{i+1}}{\longrightarrow}(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}=3\mu^{(i+1)}-4\tau^{(i+1)}+\left[\nu_{q^{(i)}}(B^{(i)})(2\nu_{q^{(i)}}(B^{(i)})-m_{q^{(i)}}(\mathcal{F}^{(i)})-1)-4\alpha_{i}\right]$
In particular
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ $=$ $3\mu^{(N)}-4\tau^{(N)}$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
$=$ $3\mu^{(N)}-4\tau^{(N)}+R_{p}(\mathcal{F},B),$
where the last equality follows from (15).
We have two possibilities: either $B\in\text{ISO}_{p}(\mathcal{F},B)$ or $B\in\text{Dic}_{p}(\mathcal{F})$. If $B\in\text{ISO}_{p}(\mathcal{F},B)$, there exists a component $D\subset\mathcal{D}$ with $D$ non-dicritical such that $\mathcal{F}^{(N)}$ has a reduced singularity at $q^{(N)}\in\mathcal{D}\cap B^{(N)}$. In this case, we have $\mu^{(N)}=\tau^{(N)}$ by Remark 4.2, and thus
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=-\tau^{(N)}+R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
If $B\in\text{Dic}_{p}(\mathcal{F})$, then there exists a component $D\subset\mathcal{D}$ with $D$ dicritical such that $B^{(N)}$ is transversal to $D$ at $q^{(N)}$, and $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$. In this case, we get
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
In both cases, the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ implies $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. $\Box$
The following example illustrates that the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ in Theorem A cannot be removed.
Example 4.3.
Let $\mathcal{F}_{m}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
where $m\in\mathbb{Z}_{>0}$. Let $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$. We have that $B$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$ such that
$GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(B)=2m(2m-1)\,\,\,\,\,\,\,\,\,\,m\geq 1,$
$\tau_{0}(B)=\left\{\begin{array}[]{cl}2&\text{if $m=1$}\\ 3m^{2}&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ 2m^{2}+2m+1&\text{ if $m>1$}\end{array}\right.$
and
$\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
A straightforward calculation reveals that $\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}<\frac{4}{3}$ for $60<m<$ 6, and
$\lim_{m\to+\infty}\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}=2.$
On the other hand, we compute $\text{DG}_{0}(\mathcal{F}_{m},B)$ for $m\geq 1$. Note that $\nu_{0}(\mathcal{F}_{m})=m+2$, and $\nu_{0}(B)=2m$, and using [19, equation 4] we get
$\alpha_{0}=\left\{\begin{array}[]{cl}1&\text{ if $m=1$}\\ m^{2}+m&\text{ if $m>1$}\end{array}\right.$
After a blow-up at $0\in\mathbb{C}^{2}$, the strict transform of $B$ is a smooth curve, and $m_{0}(\mathcal{F}_{m})=m+2$. The reduction of singularities of the pair $(\mathcal{F}_{m},B)$ is achieved in $2m$ steps, and $B\in\text{Dic}_{0}(\mathcal{F}_{m})$. Thus
$\text{DG}_{0}(\mathcal{F}_{1},B)$ $=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(1)}}(B^{(1)})(2\nu_{q^{(1)}}(B^{(1)})-m_{q^{(1)}}(\mathcal{F}^{(1)})-1)-4\alpha_{1}\right]$
$=$ $\left[2(4-3-1)-4(1)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $-5$
and for $m>1$, we have
$\text{DG}_{0}(\mathcal{F}_{m},B)$ $=$ $\sum_{j=0}^{2m-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}_{m})-1)-4\alpha_{j}\right]$
$=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F}_{2m})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(2m-1)}}(B^{(2m-1)})(2\nu_{q^{(2m-1)}}(B^{(2m-1)})-m_{q^{(2m-1)}}(\mathcal{F}^{(2m-1)}_{2m})-1)-4\alpha_{2m-1}\right]$
$=$ $\left[2m\left(2(2m)-(m+2)-1\right)-4(m^{2}+m)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $1\displaystyle 2m^{2}-10m-$ 1 .
Therefore, $\text{DG}_{0}(\mathcal{F}_{m},B)\geq 0$ if and only if $m\geq 6$. This shows that the hypothesis of Theorem A is optimal.
5. Applications
In this section, from the proof of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
Corollary 5.1.
Let $f(x,y)\in\mathbb{C}\{x,y\}$ an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
Proof.
Consider the singular holomorphic foliation $\mathcal{F}:\omega=df$. In this situation, $C\in\text{ISO}_{0}(\mathcal{F})$. To continue, we compute
$\text{DG}_{0}(\mathcal{F},C)=R_{0}(\mathcal{F},C)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)}),$
where
$R_{0}(\mathcal{F},C):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(2\nu_{q^{(j)}}(C^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
and $C^{(j+1)}$ denotes the strict transform of $C^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $C^{(N)}$ is the normalization of $C$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in C^{(N)}$ as in the proof of Theorem A.
Since $C^{(N)}$ is smooth, we have $\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})=GSV_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})$ by Remark 4.2, and from Lemma 3.1, we deduce
$\text{DG}_{0}(\mathcal{F},C)=-GSV_{0}(\mathcal{F},C)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right]$
Now, since $\mathcal{F}$ is a curve generalized non-dicritical foliation at $0\in\mathbb{C}^{2}$, and $\text{Sep}_{0}(\mathcal{F})=C$, we have $GSV_{0}(\mathcal{F},C)=0$ by [3, p. 528]. Moreover $\mu_{0}(\mathcal{F},C)=\mu(C)$ and $\tau_{0}(\mathcal{F},C)=\tau(C)$ by Remark 2.1. Hence
$\text{DG}_{0}(\mathcal{F},C)=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right].$
Finally, applying successively [19, Claim 4.4] in each step of the resolution of singularities of $C$, and given that the corresponding value $\mathcal{D}_{min}$ of [19, Claim 4.4] in the $j$-th step of the resolution of $C$ is less or equal to $\alpha_{j}$, we get $\text{DG}_{0}(\mathcal{F},C)<0$ implying
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}$
by Theorem A. ∎
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
Corollary 5.2.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. If $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$ with $\text{DG}_{p}(\mathcal{F},B_{j})<0$ for all $1\leq j\leq\ell$, then
(28) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}<\frac{4}{3}.$
Proof.
Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
$\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
Thus
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})<-\ell+1=3\ell-4\ell+(4-3)$
and
$3\left(\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-\ell+1\right)<4\left(\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})-\ell+1\right),$
so
$\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
The following example illustrates Corollary 5.2.
Example 5.3.
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=5$, $GSV_{0}(\mathcal{F},C_{1,1})=5$, $\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$, hence $\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B})}=\frac{25}{26}<\frac{4}{3}$.
We conclude the paper with a question:
Question 1: Is the inequality (28) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?