### v3

# A Dimca-Greuel type inequality for foliations

Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. We prove the following Dimca-Greuel type inequality $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$, $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$, and $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes, but it is in line with the idea developed by Wang.
Key words and phrases:
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
2020 Mathematics Subject Classification:
Primary 32S65 - 32M25
1. Introduction
Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables. Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$. The Milnor number of $C$ is
$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
and its Tjurina number is
$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
Is clear from their definitions that $\tau(C)\leq\mu(C)$.
In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [10], and [17] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}\{x,y\}/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9, p. 4], [11, Definition 3.1]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension
$\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$. If the foliation $\mathcal{F}$ is of the second type (see [9, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
We note that, in general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
On the other hand, following Camacho-Lins Neto-Sad [4, Section 4], we can consider the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition). We note that if $B=\{f(x,y)=0\}$, and $\mathcal{F}:df=0$, then $\mu_{p}(\mathcal{F},B)$ also coincides with the Milnor number $\mu(B)$ of $B$ (see Remark 2.1). Therefore, we could think to studying if the inequality
$03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0
holds. But, for the germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},0)$ defined by
(2) $\omega=y\left(13x^{7}+6y^{7}\right)dx-x\left(5y^{7}+12x^{7}\right)dy,$
and the separatrix
(3) $B:x^{13}+x^{6}y^{7}+y^{12}=0,$
we have $\mu_{0}(\mathcal{F},B)=85$ and $\tau_{0}(\mathcal{F},B)=61$. Thus
$113\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61=255-244=$ 11 .
In order to establish an inequality of the Dimca-Greuel’s type for foliations, it is necessary to add the Gómez-Mont–Seade–Verjovsky index (in short GSV-index) (see [12]) as stating our main result.
Theorem A.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
(4) $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$
where $GSV_{p}(\mathcal{F},B)$ denotes the GSV-index of $\mathcal{F}$ with respect to $B$. Moreover, the equality holds if and only if $B$ is smooth.
We will prove Theorem A in Section 3. For the foliation given in (2), and $B$ as in (3), we have $GSV_{0}(\mathcal{F},B)=-47$. Hence,
$03\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)+GSV_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61+(-47)=-36<$ 0
verifying Theorem A.
As a consequence of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17] (see Corollary 5.1).
2. Multiplicity of a foliation along a divisor of separatrices
Throughout this note $\mathcal{F}$ denotes a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$ are coprime; or by its dual vector field
$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$. Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}\{x,y\}$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, the multiplicity of $\mathcal{F}$ along $B$ at $p$ is by definition
(5) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, p. 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
(6) $\theta(t)=\begin{cases}-\frac{Q(\gamma(t))}{x^{\prime}(t)}&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{P(\gamma(t))}{y^{\prime}(t)}&\text{if $y(t)\neq 0$}.\end{cases}$
We extend the notion of multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}B$ of separatrices of $\mathcal{F}$ at $p$ in the following way:
(7) $\mu_{p}(\mathcal{F},\mathcal{B})=\left(\displaystyle\sum_{B}a_{B}\mu_{p}(\mathcal{F},B)\right)-\deg(\mathcal{B})+1.$
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
$\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
$\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$. See [12] for details.
Remark 2.1.
Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (5) and (6), we have
$\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [7, Proposition 6.2], we obtain
(8) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
Consequently when $B$ is quasihomogeneous by ([15], Satz p.123 a) and d)] and (8) or by [18, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
3. Proof of Theorem A
Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
(9) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (9):
$\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Put $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([17]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [17, Claim 4.4] we have
(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},\,\,\,\,\text{for}\,\,\,\,\nu_{p}(C)\geq 2.$
We will need the following lemma.
Lemma 3.1.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by $\sigma$ at the point $q$, then
(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Moreover, the equality holds if and only if $B$ is smooth.
Proof.
Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we have
(12) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
where
$m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behavior under blow-up of the GSV index [3, p. 30] we get
(13) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and after [7, Proposition 6.2] we have
(14) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
From [17, Equation (4)] we obtain
(15) $\tau_{p}(B)-\tau_{q}(\tilde{B})\geq\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (10). From (14), (13), and (15) we deduce
(16) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})\leq\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha.$
After (12), (13) and (16) we get
(17) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$$\geq$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
$+$ $4\alpha-\nu_{p}(B)(\nu_{p}(B)-1)$
$>$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B),$
where the last inequality follows by (10), since $\nu_{p}(B)\geq 2.$
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some convergent power series $h(x,y)\in\mathbb{C}\{x,y\}$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(18) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (14) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus, if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality. Moreover, if (11) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (17) we have that (11) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
Remark 3.2.
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(19) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (18), if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
We will now prove the main theorem of this paper.
3.1. Proof of Theorem A
Suppose that $\mathcal{F}$ is a holomorphic foliation at $p\in\mathbb{C}^{2}$, $B$ is a separatrix of $\mathcal{F}$ at $p$. First, assume that $B$ is singular, that is, $\nu_{p}(B)\geq 2$. Then, there is a sequence of blow-ups of $\mathcal{F}$ at $p$ (cf. Seidenberg [16]):
(20) $(\mathcal{F}^{(N)},B^{(N)})\longrightarrow\cdots\longrightarrow(\mathcal{F}^{(2)},B^{(2)})\longrightarrow(\mathcal{F}^{(1)},B^{(1)})\longrightarrow(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf. [11, p. 1423]). Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
Applying Lemma 3.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\longrightarrow(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}+GSV_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})\leq 3\mu^{(i+1)}-4\tau^{(i+1)}+GSV_{q^{(i+1)}}(\mathcal{F}^{(i+1)},B^{(i+1)}).$
In particular
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<3\mu^{(N)}-4\tau^{(N)}+GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})=0,$
where the last equality holds, since $B^{(N)}$ is a smooth curve.
If $B$ is a smooth curve then, by Remarks 3.2, and 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$, then the inequality (4) is an equality. Reciprocally, if (4) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by Lemma 3.1, $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<$ 0, this is a contradiction. Hence, the theorem follows. $\Box$
4. Examples
In this section, we provide two examples of holomorphic foliations and separatrices that illustrate Theorem A.
Example 4.1.
Let $\mathcal{G}_{m,n}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\eta_{m,n}=mxdy-nydx,$
where $1\leq m\leq n$ are integers. The curve $B:y^{m}-x^{n}=0$ is a separatrix of $\mathcal{G}_{m,n}$ at $0\in\mathbb{C}^{2}$ and
$GSV_{0}(\mathcal{G}_{m,n},B)$ $=$ $m+n-mn,$
$\mu_{0}(\mathcal{G}_{m,n},B)$ $=$ $1,$
$\tau_{0}(\mathcal{G}_{m,n},B)$ $=$ $1$
We get
$3\mu_{0}(\mathcal{G}_{m,n},B)-4\tau_{0}(\mathcal{G}_{m,n},B)+GSV_{0}(\mathcal{G}_{m,n},B)=m+n-mn-1=-(m-1)(n-1)\leq 0$
Thus, Theorem A is verified.
Example 4.2.
Let $\mathcal{F}_{m}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
where $m\in\mathbb{Z}_{>0}$. The curve $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We get
$GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(B)=2m(2m-1)\,\,\,\,\,\,\,\,\,\,m\geq 1,$
$\tau_{0}(B)=\left\{\begin{array}[]{cl}2&\text{if $m=1$}\\ 3m^{2}&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ 2m^{2}+2m+1&\text{ if $m>1$}\end{array}\right.$
and
$\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
A straightforward calculation reveals that
$3\mu_{0}(\mathcal{F}_{m},B)-4\tau_{0}(\mathcal{F}_{m},B)+GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}-2&\text{if $m=1$}\\ -6m&\text{ if $m>1$}.\end{array}\right.$
Therefore, Theorem A is verified.
5. Applications
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17]:
Corollary 5.1.
Let $f(x,y)\in\mathbb{C}\{x,y\}$ be an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
Proof.
It is a consequence of Remark 2.1 and applying Theorem A to the foliation $\mathcal{F}:df=0$, because $\mu_{p}(\mathcal{F},C)=\mu(C)$, $\tau_{p}(\mathcal{F},C)=\tau(C)$, and $GSV_{p}(\mathcal{F},C)=0$. ∎
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
Corollary 5.2.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(21) $03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0
Proof.
Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
$\sum_{j=1}^{\ell}\left[3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})+GSV_{p}(\mathcal{F},B_{j})\right]\leq-\ell.$
Thus
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4T_{p}(\mathcal{F},\mathcal{B})+\sum_{j=1}^{\ell}GSV_{p}(\mathcal{F},B_{j})\leq-\ell$
and, using (7) together with the GSV-index addition formula ([3, p. 29]), we obtain
$3\left(\mu_{p}(\mathcal{F},\mathcal{B})+\ell-1\right)-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})+2\sum_{i\neq k}i_{p}(B_{i},B_{k})\leq-\ell,$
so
$3\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})\leq-4\ell+3-2\sum_{i\neq k}i_{p}(B_{i},B_{k}).$
Since $\ell\geq 1$, we get
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0 .
The following example illustrates Corollary 5.2.
Example 5.3.
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=GSV_{0}(\mathcal{F},C_{1,1})=5$, so
$GSV_{0}(\mathcal{F},\mathcal{B})=GSV_{0}(\mathcal{F},C)+GSV_{0}(\mathcal{F},C_{1,1})-2i_{0}(C,C_{1,1})=5+5-2\cdot 5=0,$
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$. Hence
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})=3\cdot 25-4\cdot 26=-29<$ 0 .
We conclude the paper with a question:
Question: Is the inequality (21) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done.
Funding. This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979 and by the Spanish grant PID2019-105896GB-I00 funded by MCIN/AEI/10.13039/501100011033.
Conflict of interest. The authors of this work declare that they have no conflicts of interest.