### v4

# An upper bound for the GSV-index of a foliation

Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and let $B$ be a separatrix of $\mathcal{F}$. We prove the following upper bound $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$, where $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$, $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
Key words and phrases:
Holomorphic foliations, Gómez-Mont-Seade-Verjovsky index, Tjurina number, Multiplicity of a foliation along a divisor of separatrices.
2020 Mathematics Subject Classification:
Primary 32S65 - 32M25
1. Introduction
Let $\mathcal{F}$ be a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, and $B$ be a separatrix of $\mathcal{F}$. Several numerical invariants can be studied for the pair $(\mathcal{F},B)$, such as the Gómez-Mont–Seade–Verjovsky index (see [10]) of $\mathcal{F}$ with respect to $B$, denoted by $GSV_{p}(\mathcal{F},B)$, the multiplicity of $\mathcal{F}$ along $B$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and the Tjurina number of $\mathcal{F}$ with respect to $B:f(x,y)=0$, defined by
$\tau_{p}(\mathcal{F},B):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
In this paper, inspired by the blow-up techniques developed by Wang [14], we establish an optimal upper bound for $GSV_{P}(\mathcal{F},B)$ in terms of $\mu_{p}(\mathcal{F},B)$ and $\tau_{p}(\mathcal{F},B)$ as follows:
Theorem A.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
(1) $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$
where $GSV_{p}(\mathcal{F},B)$ denotes the GSV-index of $\mathcal{F}$ with respect to $B$. Moreover, the equality holds if and only if $B$ is smooth.
Note that if $B=\{f=0\}$ is an irreducible plane curve germ, then by applying Theorem A to the foliation $\mathcal{F}:df=0$, we obtain the bound stated in [6, Question 4.2] and proved simultaneously in [1], [8] and [14], since $\mu_{p}(\mathcal{F},B)=\mu(B)$, $\tau_{p}(\mathcal{F},B)=\tau(B)$, and $GSV_{p}(\mathcal{F},B)=0$ in this case. Here, $\mu(B)$ and $\tau(B)$ denote the classical Milnor and Tjurina numbers of $B$. A complete answer to this question was given by Almirón [2]. Moreover, the author proposed a broader perspective of the study of the difference between Milnor and Tjurina numbers for isolated complete intersection singularities.
2. Multiplicity of a foliation along a divisor of separatrices
Throughout this note $\mathcal{F}$ denotes a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$ are coprime; or by its dual vector field
$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$. Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}\{x,y\}$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, the multiplicity of $\mathcal{F}$ along $B$ at $p$ is by definition
(2) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, p. 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
(3) $\theta(t)=\begin{cases}-\frac{Q(\gamma(t))}{x^{\prime}(t)}&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{P(\gamma(t))}{y^{\prime}(t)}&\text{if $y(t)\neq 0$}.\end{cases}$
We extend the notion of multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}B$ of separatrices of $\mathcal{F}$ at $p$ in the following way:
(4) $\mu_{p}(\mathcal{F},\mathcal{B})=\left(\displaystyle\sum_{B}a_{B}\mu_{p}(\mathcal{F},B)\right)-\deg(\mathcal{B})+1.$
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [11, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
$\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
$\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$. See [10] for details.
Remark 2.1.
Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (2) and (3), we have
$\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [7, Proposition 6.2], we obtain
(5) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
Consequently when $B$ is quasihomogeneous by ([12], Satz p.123 a) and d)] and (5) or by [15, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
3. Proof of Theorem A
Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
(6) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (6):
$\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Put $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([14]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [14, Claim 4.4] we have
(7) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},\,\,\,\,\text{for}\,\,\,\,\nu_{p}(C)\geq 2.$
We will need the following lemma.
Lemma 3.1.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by $\sigma$ at the point $q$, then
(8) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Moreover, the equality holds if and only if $B$ is smooth.
Proof.
Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we have
(9) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
where
$m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behavior under blow-up of the GSV index [3, p. 30] we get
(10) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and after [7, Proposition 6.2] we have
(11) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
From [14, Equation (4)] we obtain
(12) $\tau_{p}(B)-\tau_{q}(\tilde{B})\geq\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (7). From (11), (10), and (12) we deduce
(13) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})\leq\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha.$
After (9), (10) and (13) we get
(14) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$$\geq$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
$+$ $4\alpha-\nu_{p}(B)(\nu_{p}(B)-1)$
$>$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B),$
where the last inequality follows by (7), since $\nu_{p}(B)\geq 2.$
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some convergent power series $h(x,y)\in\mathbb{C}\{x,y\}$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(15) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (11) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus, if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (8) becomes an equality. Moreover, if (8) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by (7) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (14) we have that (8) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
Remark 3.2.
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(16) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (15), if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
We will now prove the main theorem of this paper.
3.1. Proof of Theorem A
Suppose that $\mathcal{F}$ is a holomorphic foliation at $p\in\mathbb{C}^{2}$, $B$ is a separatrix of $\mathcal{F}$ at $p$. First, assume that $B$ is singular, that is, $\nu_{p}(B)\geq 2$. Then, there is a sequence of blow-ups of $\mathcal{F}$ at $p$ (cf. Seidenberg [13]):
(17) $(\mathcal{F}^{(N)},B^{(N)})\longrightarrow\cdots\longrightarrow(\mathcal{F}^{(2)},B^{(2)})\longrightarrow(\mathcal{F}^{(1)},B^{(1)})\longrightarrow(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf. [9, p. 1423]). Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
Applying Lemma 3.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\longrightarrow(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}+GSV_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})\leq 3\mu^{(i+1)}-4\tau^{(i+1)}+GSV_{q^{(i+1)}}(\mathcal{F}^{(i+1)},B^{(i+1)}).$
In particular
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<3\mu^{(N)}-4\tau^{(N)}+GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})=0,$
where the last equality holds, since $B^{(N)}$ is a smooth curve.
If $B$ is a smooth curve then, by Remarks 3.2, and 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$, then the inequality (1) is an equality. Reciprocally, if (1) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by Lemma 3.1, $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<$ 0, this is a contradiction. Hence, the theorem follows. $\Box$
4. Examples
In this section, we provide two examples of holomorphic foliations and separatrices that illustrate Theorem A.
Example 4.1.
Let $\mathcal{G}_{m,n}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\eta_{m,n}=mxdy-nydx,$
where $1\leq m\leq n$ are integers. The curve $B:y^{m}-x^{n}=0$ is a separatrix of $\mathcal{G}_{m,n}$ at $0\in\mathbb{C}^{2}$ and
$GSV_{0}(\mathcal{G}_{m,n},B)$ $=$ $m+n-mn,$
$\mu_{0}(\mathcal{G}_{m,n},B)$ $=$ $1,$
$\tau_{0}(\mathcal{G}_{m,n},B)$ $=$ $1$
We get
$3\mu_{0}(\mathcal{G}_{m,n},B)-4\tau_{0}(\mathcal{G}_{m,n},B)+GSV_{0}(\mathcal{G}_{m,n},B)=m+n-mn-1=-(m-1)(n-1)\leq 0$
Thus, Theorem A is verified.
Example 4.2.
Let $\mathcal{F}_{m}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
where $m\in\mathbb{Z}_{>0}$. The curve $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We get
$GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(B)=2m(2m-1)\,\,\,\,\,\,\,\,\,\,m\geq 1,$
$\tau_{0}(B)=\left\{\begin{array}[]{cl}2&\text{if $m=1$}\\ 3m^{2}&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ 2m^{2}+2m+1&\text{ if $m>1$}\end{array}\right.$
and
$\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
A straightforward calculation reveals that
$3\mu_{0}(\mathcal{F}_{m},B)-4\tau_{0}(\mathcal{F}_{m},B)+GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}-2&\text{if $m=1$}\\ -6m&\text{ if $m>1$}.\end{array}\right.$
Therefore, Theorem A is verified.
5. Applications
We define the sum $T_{p}(\mathcal{F},C):=\sum_{B\subset C}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality for foliations:
Corollary 5.1.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(18) $GSV_{p}(\mathcal{F},\mathcal{B})<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})$
Proof.
Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
$\sum_{j=1}^{\ell}\left[3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})+GSV_{p}(\mathcal{F},B_{j})\right]\leq-\ell.$
Thus
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4T_{p}(\mathcal{F},\mathcal{B})+\sum_{j=1}^{\ell}GSV_{p}(\mathcal{F},B_{j})\leq-\ell$
and, using (4) together with the GSV-index addition formula ([3, p. 29]), we obtain
$3\left(\mu_{p}(\mathcal{F},\mathcal{B})+\ell-1\right)-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})+2\sum_{i\neq k}i_{p}(B_{i},B_{k})\leq-\ell,$
so
$3\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})\leq-4\ell+3-2\sum_{i\neq k}i_{p}(B_{i},B_{k}).$
Since $\ell\geq 1$, we get
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0 .
The following example illustrates Corollary 5.1.
Example 5.2.
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=GSV_{0}(\mathcal{F},C_{1,1})=5$, so
$GSV_{0}(\mathcal{F},\mathcal{B})=GSV_{0}(\mathcal{F},C)+GSV_{0}(\mathcal{F},C_{1,1})-2i_{0}(C,C_{1,1})=5+5-2\cdot 5=0,$
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$. Hence
$GSV_{p}(\mathcal{F},\mathcal{B})=0<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})=29.$
We conclude the paper with a question:
Question: Is the inequality (18) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done.
Funding. This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979 and by the Spanish grant PID2019-105896GB-I00 funded by MCIN/AEI/10.13039/501100011033.
Conflict of interest. The authors of this work declare that they have no conflicts of interest.